OpenStudy (anonymous):

How would I solve cos (sin x) = 1?

OpenStudy (anonymous):

Focus on \(\cos u=1\) first. What could make \(\cos\) yield \(1\)? Well, how about \(u=2\pi k\) for integer \(k\) -- right?

OpenStudy (anonymous):

Now the question is, when does \(\sin x=2\pi k\)?

OpenStudy (anonymous):

Well, observe that \(\sin x\in[-1,1]\) -- and \(-1\le 2\pi k\le 1\) only for \(k=0\), correct?

OpenStudy (anonymous):

2pik should belong to [-1,1]

OpenStudy (anonymous):

... so it boils down to \(\sin x=0\) -- can you solve this?

OpenStudy (anonymous):

I'm still a bit confused :/

OpenStudy (anonymous):

@RPguy in solving \(\cos(\sin x)=0\), we're first interested in what \(\sin x\) has to be to make \(\cos(\sin x)=0\). Once we figure that out, we have to figure out what \(x\) has to be to yield those possible values of \(\sin x\).

OpenStudy (anonymous):

Okay, that makes more sense

OpenStudy (anonymous):

@RPguy so we observe that to make \(\cos(\sin x)=0\) we'd need \(\sin x\) to be some integer multiple of \(2\pi\), right? convince yourself of this by looking at the unit circle. Now, what integer multiples in particular are possible here?

OpenStudy (anonymous):

Think of the "range" of \(\sin x\) -- it yields numbers between \(-1\) and \(1\), inclusive, correct? What integer multiples of \(2\pi\) lie in this range? Once you determine that, figure out what values of \(x\) could yield those values of \(\sin x\) :-)

OpenStudy (anonymous):

Okay, so basically \[\cos (\sin 0)=1\]

OpenStudy (anonymous):

@RPguy the only integer multiple in our desired range is \(0\), hence our problem reduces to solving \(\sin x = 0\). Here, however, we have more flexibility; we find \(x\) can be any integer multiple of \(\pi\). Hence \(x=\pi n\) for integer \(n\).

OpenStudy (anonymous):

Why wouldn't it be \[n2\pi\]

OpenStudy (jdoe0001):

it'd depend on the value of "n" I guess otherwise sin(x) is 0 at 0 + \(\bf k\pi\)

OpenStudy (jdoe0001):

hmm, oh, cosine has to be hmmm 1

OpenStudy (jdoe0001):

yeah, \(n2\pi\) looks good

OpenStudy (anonymous):

@jdoe0001 Okay thanks :)

OpenStudy (anonymous):

No, \(2\pi k\) misses all the odd multiples and is an incomplete solution.